OpenStudy (anonymous):

4 over 3x + 2 over 3 = 6

OpenStudy (anonymous):

\[\frac{4}{3x}+\frac{2}{3}=6 \implies \frac{4}{3x}=\frac{16}{3} \implies 48x=12 \implies x=\frac{1}{4}\]